OpenStudy (anonymous):

Suppose a bank loans $7,200 at an interest rate of 4% compounded annually. Find the total amount owed on the loan after 4 years. A. $7787.52 B. $8422.98 C. $27,659.52 D. $29,952.00

OpenStudy (anonymous):

@Juicstice

OpenStudy (amistre64):

id just use the formula

OpenStudy (amistre64):

\[A=P(1+r/n)^{nt}\]

OpenStudy (amistre64):

since it compounds once a year, n=1 and the rest is pretty much self evident

OpenStudy (dumbcow):

P = 7200 r = .04 t =4

OpenStudy (anonymous):

Here's an example of how to use the formula that aministre said... A=P(1+r/n)^nt [Use the compound interest formula] A=The amount of money in the savings account P=The principle r=The interest rate converted to a decimal t=The time that the money is in the account n=The number of times the money is compounded per year . Write each answer to the following questions in a complete sentence. a. If the principle amount is $800 at 2.4% interest compounded monthly for 5 years the total amount will be $901.89. A = 800 [ 1 + ((.024)/(12)) ]^((12)(5)) A =901.89 . b. If a person wishes to have $1500 in 4 years and it is compounded quarterly at an interest rate of 3.5%, she must invest $1304.83. A=P(1+(r/n))^(nt) [Solve for P] P=(1/[1+(r/n))^(nt)] [Plug-in the values] P = (1500) / {[ 1 + ((.035)/(4)) ]^((4)(4))} P=1304.8316 .

OpenStudy (anonymous):

*amistre

OpenStudy (anonymous):

can someone telll me how I do the formula...Im not getting this!

OpenStudy (anonymous):

@amistre64

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ p=\textit{ original amount}\to 7,200\\ n=\textit{compounding cycle per year, annually means once a year}\to 1\\ r=rate\to 4\%\to \frac{4}{100}\to 0.04\\ t=years\to 4 \\ \quad \\ A=7,200\left(1+\frac{0.04}{1}\right)^{1\cdot 4}\)